# Ke Xua,b, Fen-Fen Yanga,b, Chenggui Yuanc

### v1

@@ Abstract
In this work, we establish the existence and uniqueness of solutions to McKean-Vlasov stochastic differential equations (SDEs) driven by Lévy processes with common noise on an infinite time horizon, by means of a contraction mapping principle in the space of probability measures. In addition, we analyse the propagation of chaos for Lévy-driven McKean-Vlasov SDEs in the presence of common noise.
Keywords: McKean-Vlasov equations; Lévy processes; Common noise; Moment estimates; Propagation of chaos
@@ 1 Introduction
for $t\geq 0$, with initial value $X(0)=X_{0}$ and its distribution $\mu=\mathcal{L}(X_{0})$. Here $\mathcal{L}^{1}(X_{t})$ denotes the marginal law of $X$ at time $t$, $W=(W_{t})_{t\geq 0}$ is a $d$-dimensional standard Brownian motion. Denote $\mathbb{R}_{0}^{d}:=\mathbb{R}^{d}\setminus\{0\}$ and let $\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d})$ be the Borel $\sigma$-algebra on $\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}$.
@@
where $\widetilde{N}$ is the compensated Poisson random measure associated with the Poisson random measure $N$ on $(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d},\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}))$ with intensity measure $\nu(\mathrm{d}z)\mathrm{d}s$. The compensated Poisson random measure is given by
$\widetilde{N}(\mathrm{d}t,\mathrm{d}z):=N(\mathrm{d}t,\mathrm{d}z)-\nu(\mathrm{d}z)\mathrm{d}t.$
Explicitly, for $t>0$ and $A\in\mathcal{B}(\mathbb{R}_{0}^{d})$,
@@
Let $|\cdot|$ denote the Euclidean norm on $\mathbb{R}^{d}$. Let $(\mathcal{F}^{0}_{t})_{t\geq 0}$, $(\mathcal{F}^{1}_{t})_{t\geq 0}$ be two complete filtered probability spaces satisfying the usual condition. Let $(W_{t}^{0})_{t\geq 0}$ and $(W_{t})_{t\geq 0}$ be independent $d$-dimensional Brownian motions defined on $(\Omega^{0},\mathcal{F}^{0},\mathbb{P}^{0})$ and $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, respectively. We construct the product probability space $(\Omega,\mathcal{F},\mathbb{P})$ by setting
@@
which can be interpreted as a mapping from $\Omega^{0}$ into $\mathcal{P}(\mathbb{R}^{d})$. This object represents the conditional distribution of $\xi$ given $\mathcal{F}^{0}$. In particular, if the initial condition $X(0)$ in (1.1) is assumed to be defined on $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, then $\mathcal{L}^{1}(X(t))$ provides a version of the conditional law of $X(t)$ given $W^{0}$. Henceforth, we shall restrict our attention to initial conditions living on $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$.
@@
The existence and uniqueness theory for strong solutions of McKean-Vlasov SDEs under Lipschitz and linear growth conditions (with respect to both the state variable and the measure) is well-established (see, e.g., [34], [38]). Wang [42] establishes well-posedness via Picard iteration method resolving pathwise intractability in the Brownian-driven case. In the context of delay McKean-Vlasov SDEs with super-linear drift, Bao et al. [5] propose a Milstein-based antithetic MLMC framework attaining optimal complexity without Lévy area simulation under Brownian noise. Liu et al. [24] establish strong and weak well-posedness of McKean-Vlasov SDEs under locally monotone coefficients in state and distribution variables, and derived a large deviation principle with applications to granular media, plasma, kinetic, porous media, and Navier-Stokes equations.
@@
Compared with Brownian-driven McKean-Vlasov SDEs, the McKean-Vlasov SDEs driven by Lévy processes have been much less studied. Bao and Wang [7] establish existence, uniqueness, and multiplicity of stationary distributions for Lévy-driven McKean-Vlasov SDEs. Song et al. [30] develop large and moderate deviation principles, along with explicit rate functions, for Lévy-driven McKean-Vlasov SDEs using a weak-convergence approach. In [26], Jourdain et al. investigate the existence, uniqueness, and particle approximations for McKean-Vlasov SDEs driven by Lévy processes. More recently, Mehri et al. [35] and Neelima et al. [37] address the well-posedness and propagation of chaos results for McKean-Vlasov SDEs with delay driven by Lévy noise.
@@
The Brownian-driven McKean-Vlasov SDEs with common noise have been applied to various fields including finance, physics, biology, mean-field games, and machine learning. The fundamental theory of the McKean-Vlasov SDEs with common noise is well established. One can refer to Carmona and Delarue [13], Kumar et al. [27] for the well-posedness and conditional propagation of chaos, Hammersley et al. [25] for the weak well-posedness, Maillet [36], Bao and Wang [8] for the long-time behavior. Yuan et al. [16] design delay feedback control strategies for stabilizing McKean-Vlasov SDEs with common noise, proving the global solution existence and uniqueness. Shao and Wei [39] study the well-posedness and conditional propagation of chaos for McKean-Vlasov SDEs with common noise, establishing explicit convergence rates.
@@
An epidemic outbreak occurs when a contagious disease spreads rapidly through a population due to increased transmission rates. In such circumstances, it is reasonable to incorporate a jump process to capture sudden surges in infection numbers triggered by super-spreader events. Based on this perspective, the evolution of the infected population can be modelled by a jump diffusion. Moreover, when global mobility is high, localized outbreaks can quickly propagate across regions, resulting in a pandemic (i.e., a worldwide health crisis in which many areas experience simultaneous infection waves). This observation shows that epidemic dynamics are subject to system-wide randomness (e.g., environmental factors, global travel) that impact all communities. In deep learning, training dynamics can be modeled by Lévy-driven McKean-Vlasov SDEs, where parameter updates follow distribution-dependent stochastic evolution with heavy-tailed noise. Mini-batch gradients and restarts induce jump-like perturbations, while the mean-field term captures interactions among parameters in wide-network limits. In federated or distributed settings, common randomness (e.g., global schedules, network delays) produces correlated dynamics across learners. This framework helps explain heavy-tailed gradient noise, escape from saddles via jumps, and inspires neural SDE architectures that combine diffusion with jumps for expressive modeling. The above considerations motivate the study of Lévy-driven McKean-Vlasov SDEs with common noise [[9],[10]]. In recent years, significant advances have been achieved for Lévy-driven McKean-Vlasov SDEs under common noise, covering topics such as well-posedness, deep learning approaches, optimal stopping problems, optimal and impulse control, as well as stochastic maximum principles; see, for instance, [[1], [2], [3], [9], [10], [22], [23]] for further references.
@@
This paper investigates the infinite time horizon approximation for Lévy-driven McKean-Vlasov SDEs with common noise. The primary contribution of this study lies in addressing a research gap by extending previous work on McKean-Vlasov SDE with Lévy-driven process, where the drift and diffusion coefficients are not globally Lipschitz continuous. Tran et al. [41] propose a tamed-adaptive Euler-Maruyama scheme for Lévy-driven McKean-Vlasov SDEs with superlinear and non-globally Lipschitz coefficients, proving strong convergence on finite and infinite horizons. In particular, this study builds upon the results of Tran et al. [41], introducing the effect of common noise-an important factor that has not been examined in the context of such SDEs. The incorporation of common noise enhances the model, providing a more realistic framework for situations where noise simultaneously impacts all components, a consideration that has often been neglected in prior research.
@@
    We extend the results of Tran et al. [41] by establishing the boundedness of high-order moments of the solutions and demonstrating the propagation of chaos in the presence of common noise.
@@ 2 Well-posedness of Solutions
(A3) There exist constants $L>0$ and $\ell\geq 0$ such that
$|b(x,\mu)-b(\overline{x},\overline{\mu})|\leq L(1+|x|^{\ell}+|\overline{x}|^{\ell})\left(|x-\overline{x}|+\mathcal{W}_{2}(\mu,\overline{\mu})\right),$
@@
for any $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, where $\gamma=\gamma_{1}+\gamma_{2}$ represents the dissipation rate.
@@
To make the study easy to follow, we divide the proof into the following steps.

### v2

@@ Abstract
In this work, we establish the existence and uniqueness of solutions to McKean-Vlasov stochastic differential equations (SDEs) driven by Lévy processes with common noise, by means of a contraction mapping principle in the space of probability measures. In addition, on the infinite time horizon we analyse the propagation of chaos for McKean-Vlasov SDEs driven by Lévy processes in the presence of common noise, and provide numerical simulations to validate the corresponding model.
Keywords: McKean-Vlasov equations; Lévy processes; Common noise; Propagation of chaos; Numerical simulation
@@ 1 Introduction
for $t\geq 0$, with initial value $X_{0}=x_{0}$. Here $\mathcal{L}^{1}(X(t))$ provides a version of the conditional law of $X(t)$ given $W^{0}$, $W=(W_{t})_{t\geq 0}$ is a $d$-dimensional standard Brownian motion. Denote $\mathbb{R}_{0}^{d}:=\mathbb{R}^{d}\setminus\{0\}$ and let $\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d})$ be the Borel $\sigma$-algebra on $\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}$, and
@@
where $\widetilde{N}(\mathrm{d}t,\mathrm{d}z):=N(\mathrm{d}t,\mathrm{d}z)-\nu(\mathrm{d}z)\mathrm{d}t$ is the compensated Poisson random measure associated with the Poisson random measure $N$ on $(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d},\mathcal{B}(\mathbb{R}_{+}\times\mathbb{R}_{0}^{d}))$ with intensity measure $\nu(\mathrm{d}z)\mathrm{d}s$. Explicitly, for $t>0$ and $A\in\mathcal{B}(\mathbb{R}_{0}^{d})$,
@@
Let $|\cdot|$ denote the Euclidean norm on $\mathbb{R}^{d}$, and let $(\mathcal{F}^{0}_{t})_{t\geq 0}$, $(\mathcal{F}^{1}_{t})_{t\geq 0}$ be two complete filtered probability spaces satisfying the usual condition. Let $(W_{t}^{0})_{t\geq 0}$ and $(W_{t})_{t\geq 0}$ be independent $d$-dimensional Brownian motions defined on $(\Omega^{0},\mathcal{F}^{0},\mathbb{P}^{0})$ and $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, respectively. We construct the product probability space $(\Omega,\mathcal{F},\mathbb{P})$ by setting
@@
which represents the conditional distribution of $\xi$ given $\mathcal{F}^{0}$. In particular, if the initial condition $X(0)$ in (1.1) is assumed to be defined on $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$, then $\mathcal{L}^{1}(X(t))$ provides a version of the conditional law of $X(t)$ given $W^{0}$. Henceforth, we shall restrict our attention to initial conditions living on $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1})$.
@@
The existence and uniqueness theory for strong solutions of McKean-Vlasov SDEs under Lipschitz and linear growth conditions (with respect to both the state variable and the measure) is well-established (see, e.g., [25], [29]). Wang [33] establishes well-posedness via the Picard iteration method, resolving the pathwise intractability in the case of SDEs driven by Brownian motion. In the context of delay McKean-Vlasov SDEs with super-linear drift, Bao et al. [4] propose a Milstein-based antithetic MLMC framework attaining optimal complexity without Lévy area simulation under Brownian noise. Liu et al. [19] establish concentration inequalities, exponential convergence in the Wasserstein metric $\mathcal{W}_{1}$, and uniform-in-time propagation of chaos for McKean-Vlasov systems via reflection coupling and new cost functions, with applications to multi-well confinement and bounded interaction potentials.
@@
Compared with McKean-Vlasov SDEs driven by Brownian motion, the McKean-Vlasov SDEs driven by Lévy processes have been much less studied. Bao and Wang [5] establish existence, uniqueness, and multiplicity of stationary distributions for McKean-Vlasov SDEs driven by Lévy processes. Song et al. [23] develop large and moderate deviation principles, along with explicit rate functions, for McKean-Vlasov SDEs driven by Lévy processes using a weak-convergence approach. In [21], Jourdain et al. investigate the existence, uniqueness, and particle approximations for McKean-Vlasov SDEs driven by Lévy processes. More recently, Mehri et al. [26] and Neelima et al. [28] address the well-posedness and propagation of chaos results for McKean-Vlasov SDEs with delay driven by Lévy noise. Bao et al. [7] establish strong well-posedness of McKean-Vlasov SDEs driven by Lévy processes under weak monotonicity and coercivity, prove both weak and strong propagation of chaos, and extend the results to systems with common noise.
@@
The McKean-Vlasov SDEs driven by Brownian motion with common noise have been applied to various fields including finance, physics, biology, mean-field games, and machine learning. The fundamental theory of the McKean-Vlasov SDEs with common noise is well established. One can refer to Carmona and Delarue [10], Kumar et al. [22] for the well-posedness and conditional propagation of chaos, Hammersley et al. [20] for the weak well-posedness, Maillet [27], Bao and Wang [6] for the long-time behavior. Yuan et al. [11] design delay feedback control strategies for stabilizing McKean-Vlasov SDEs with common noise, proving the global solution existence and uniqueness. Shao and Wei [30] study the well-posedness and conditional propagation of chaos for McKean-Vlasov SDEs with common noise, establishing explicit convergence rates.
@@
In applications, random jump shocks naturally arise in a variety of systems. For instance, an epidemic outbreak occurs when a contagious disease spreads rapidly through a population due to increased transmission rates. Super-spreader events may cause sudden surges in infection numbers, which can be modelled by a jump component in a diffusion-type dynamics for the infected population. When global mobility is high, localized outbreaks can quickly propagate across regions, resulting in a pandemic, and the epidemic dynamics are then subject to system-wide sources of randomness (e.g., environmental factors, global travel) that impact all communities. Similar large-scale effects appear in economics and finance. In systemic risk modelling, the balance sheet of each bank is influenced not only by its idiosyncratic shocks but also by macroeconomic factors and market-wide events. This can be described by McKean-Vlasov SDEs with common noise and jumps, where the mean-field term captures peer effects and the jump component models sudden losses or liquidity dry-ups; see the interbank lending model studied later. Bistable dynamics with random environment, as in a Ginzburg-Landau type equation, provide a prototype for metastable transitions under heavy-tailed shocks, and will serve as our first test model. In conditional mean-field control problems, agents interact through conditional expectations of their states while being exposed to common noise representing shared information or policy shocks. Linear-quadratic structures then lead to tractable feedback controls and motivate the LQ test models considered in Section 5. These examples motivate the study of McKean-Vlasov SDEs driven by Lévy processes with common noise; see, for instance, [8, 9]. In recent years, significant advances have been achieved for McKean-Vlasov SDEs driven by Lévy processes under common noise, covering topics such as well-posedness, deep learning approaches, optimal stopping problems, optimal and impulse control, as well as stochastic maximum principles; see, among others, [1, 2, 3, 8, 9, 13, 15, 17, 18, 24] for further references.
@@
This paper investigates the infinite time horizon approximation for McKean-Vlasov SDEs driven by Lévy processes with common noise. The primary contribution of this study lies in addressing a research gap by extending previous work to McKean-Vlasov SDE driven by Lévy processes, where the drift and diffusion coefficients are not globally Lipschitz continuous. Tran et al. [32] propose a tamed-adaptive Euler-Maruyama scheme for McKean-Vlasov SDEs driven by Lévy processes with superlinear coefficients, proving strong convergence on finite and infinite horizons. In particular, this study builds upon the results of Tran et al. [32] by introducing the effect of common noise. This factor has not been examined in the context of such SDEs. The incorporation of common noise enhances the model, providing a more realistic framework for situations where noise simultaneously impacts all components, a consideration that has often been neglected in prior research.
@@
    We extend the results of Tran et al. [32] to establish the boundedness of high-order moments of the solutions and demonstrate the propagation of chaos in the presence of common noise.
@@ 2 Well-posedness of Solutions
(A3) There exist constants $\tilde{L}>0$ and $\ell>0$ such that
$|b(x,\mu)-b(\overline{x},\overline{\mu})|\leq\tilde{L}(1+|x|^{\ell}+|\overline{x}|^{\ell})\left(|x-\overline{x}|+\mathcal{W}_{2}(\mu,\overline{\mu})\right),$
@@
for any $x\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
@@
Remark 2.1.
Similar with Tran et al. [32]. It follows from Assumption (A6) that, for the even integer $p_{0}\in[2,+\infty)$ given in (A4) and the positive constant $L_{3}$ given in (A5), there exist constants $\gamma_{1}\in\mathbb{R}$, $\gamma_{2}\geq 0$, and $\eta\geq 0$ such that
$\langle x,b(x,\mu)\rangle+\frac{p-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})+\frac{1}{2L_{3}}|c(x,\mu)|^{2}\int_{\mathbb{R}^{d}_{0}}|z|\left(\left(1+L_{3}|z|\right)^{p-1}-1\right)\nu({{\rm{d}}}z)$
$\leq\gamma_{1}|x|^{2}+\gamma_{2}\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\eta,$
Example 1.
To illustrate that the aforementioned hypotheses are valid, we consider the following example: a variant of the Ginzburg-Landau equation. Specifically, consider
[…]
where $\mu_{t}=\mathcal{L}^{1}(X_{t})$ denotes the conditional distribution of $X_{t}$, $m(\mu_{t})=\int y\,\mu_{t}({\rm d}y)$ is its first moment, $W$ is a Brownian motion, $W^{0}$ a common Brownian motion, and $Z$ a pure-jump Lévy process with Lévy measure $\nu({\rm d}z)$. For the jump part we choose a tempered stable Lévy measure on $\mathbb{R}_{0}$:
$\nu(\mathrm{d}z)=C_{\alpha,\lambda}\,|z|^{-1-\alpha}e^{-\lambda|z|}\,\mathrm{d}z,\qquad\alpha\in(0,1),\;\lambda>0,$
where $C_{\alpha,\lambda}$ is determined by the normalization condition ensuring that $\nu$ is a valid Lévy measure, i.e.
$C_{\alpha,\lambda}=\frac{\lambda^{\alpha}}{2\Gamma(-\alpha)},$
with $\Gamma(\cdot)$ denoting the Gamma function. Note that $\Gamma(s)=\int_{0}^{\infty}t^{s-1}e^{-t}\,\mathrm{d}t$ is the Gamma function. By setting $s=q-\alpha$, we have
$\int_{\mathbb{R}_{0}}|z|^{q}\,\nu(\mathrm{d}z)=C_{\alpha,\lambda}\int_{\mathbb{R}_{0}}|z|^{q-1-\alpha}e^{-\lambda|z|}\,\mathrm{d}z=2C_{\alpha,\lambda}\,\lambda^{\alpha-q}\,\Gamma(q-\alpha),\qquad q>\alpha.$
For $q=2$, we obtain
$\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)=2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha),\qquad 0<\alpha<1.$
Next we verify that the above equation satisfies assumptions (A1)-(A7).
(1) Since $b(x,\mu)=0.1(x-x^{3}+m(\mu))$, we have
$2\langle x,b(x,\mu)\rangle$ $=2x\cdot 0.1\big(x-x^{3}+m(\mu)\big)=0.2\big(x^{2}-x^{4}+xm(\mu)\big).$
Moreover,
$|\sigma(x,\mu)|^{2}=(0.1x)^{2}=0.01x^{2},\qquad|\sigma^{0}(x,\mu)|^{2}=(0.05x)^{2}=0.0025x^{2}.$
Also we have $|c(x,\mu)|^{2}\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)=\sin^{2}(x)\cdot 2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)\leq 2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$ and $m(\mu)^{2}\leq\mathcal{W}_{2}^{2}(\mu,\delta_{0}).$ Thus, we have
$2\langle x,b(x,\mu)\rangle+|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2}+|c(x,\mu)|^{2}\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)$
$\quad=0.2(x^{2}-x^{4}+xm(\mu))+0.0125x^{2}+\sin^{2}(x)\cdot 2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$
$\quad\leq 0.2x^{2}+0.0125x^{2}+0.2|x||m(\mu)|+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$
$\quad\leq(0.2125+0.1)x^{2}+0.1\mathcal{W}_{2}^{2}(\mu,\delta_{0})+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$
$\quad=0.3125x^{2}+0.1\mathcal{W}_{2}^{2}(\mu,\delta_{0})+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha).$
Letting $L=\max\{0.3125,0.1,2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)\}+1$, we show that the Assumption (A1) holds.
(2) Note that
$|\sigma(x)-\sigma(\overline{x})|^{2}=(0.1)^{2}|x-\overline{x}|^{2}=0.01|x-\overline{x}|^{2},$
$|\sigma^{0}(x)-\sigma^{0}(\overline{x})|^{2}=(0.05)^{2}|x-\overline{x}|^{2}=0.0025|x-\overline{x}|^{2}.$
For the jump coefficient, use $|\sin x-\sin y|\leq|x-y|$, so
$|c(x)-c(\overline{x})|^{2}\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)\leq 2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)|x-\overline{x}|^{2}.$
Hence,
$2\langle x-\overline{x},b(x,\mu)-b(\overline{x},\overline{\mu})\rangle$
$\quad=2(x-\overline{x})\cdot 0.1\big((x-x^{3})-(\overline{x}-\overline{x}^{3})+m(\mu)-m(\overline{\mu})\big)$
$\quad=0.2(x-\overline{x})\big((x-x^{3})-(\overline{x}-\overline{x}^{3})\big)+0.2(x-\overline{x})(m(\mu)-m(\overline{\mu})).$
Using the algebraic identity $(x-x^{3})-(\overline{x}-\overline{x}^{3})=(x-\overline{x})\big(1-(x^{2}+x\overline{x}+\overline{x}^{2})\big),$ we have
$0.2(x-\overline{x})\big((x-x^{3})-(\overline{x}-\overline{x}^{3})\big)=0.2|x-\overline{x}|^{2}\big(1-(x^{2}+x\overline{x}+\overline{x}^{2})\big)\leq 0.2|x-\overline{x}|^{2}.$
Moreover,
$0.2(x-\overline{x})(m(\mu)-m(\overline{\mu}))\leq 0.1|x-\overline{x}|^{2}+0.1|m(\mu)-m(\overline{\mu})|^{2}.$
Combining the above bounds gives
$\leq\big(0.2+0.1+0.01+0.0025+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)\big)|x-\overline{x}|^{2}+0.1|m(\mu)-m(\overline{\mu})|^{2}$
$=\Big(0.3125+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)\Big)|x-\overline{x}|^{2}+0.1|m(\mu)-m(\overline{\mu})|^{2}.$
Since $|m(\mu)-m(\overline{\mu})|\leq\mathcal{W}_{2}(\mu,\overline{\mu})$, we may choose
$L_{1}:=0.3125+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha),\qquad L_{2}:=0.1.$
Hence Assumption (A2) holds with these constants.
(3) Due to
$|b(x,\mu)-b(\overline{x},\overline{\mu})|$ $=0.1\big|(x-x^{3})-(\overline{x}-\overline{x}^{3})+m(\mu)-m(\overline{\mu})\big|$
$\leq 0.1\big|(x-x^{3})-(\overline{x}-\overline{x}^{3})\big|+0.1|m(\mu)-m(\overline{\mu})|.$
Applying the mean-value theorem: there exists $\xi$ between $x$ and $\overline{x}$ such that
$|(x-x^{3})-(\overline{x}-\overline{x}^{3})|=|x-\overline{x}||1-3\xi^{2}|\leq|x-\overline{x}|(1+3\xi^{2})\leq|x-\overline{x}|(1+3(|x|^{2}+|\overline{x}|^{2})).$
Thus
$|b(x,\mu)-b(\overline{x},\overline{\mu})|\leq 0.1\big(1+3(|x|^{2}+|\overline{x}|^{2})\big)|x-\overline{x}|+0.1\mathcal{W}_{2}(\mu,\overline{\mu}).$
This implies that Assumption (A3) holds with $\ell=2$.
(4) Note that $\Gamma(s,\lambda)=\int_{\lambda}^{\infty}t^{s-1}e^{-t}\,\mathrm{d}t$ is the upper incomplete Gamma function. When $2p_{0}>\alpha$, we have
$\int_{|z|>1}|z|^{2p_{0}}\,\nu(\mathrm{d}z)$ $=2C_{\alpha,\lambda}\int_{1}^{\infty}r^{2p_{0}-1-\alpha}e^{-\lambda r}\,\mathrm{d}r$
$=2C_{\alpha,\lambda}\,\lambda^{\alpha-2p_{0}}\,\Gamma(2p_{0}-\alpha,\lambda).$
Similarly, due to $\gamma(s,\lambda)=\int_{0}^{\lambda}t^{s-1}e^{-t}\,\mathrm{d}t$ is the lower incomplete Gamma function. When $\alpha<1$, we have
$\int_{0<|z|\leq 1}|z|\,\nu(\mathrm{d}z)$ $=2C_{\alpha,\lambda}\int_{0}^{1}r^{-\alpha}e^{-\lambda r}\,\mathrm{d}r$
$=2C_{\alpha,\lambda}\,\lambda^{\alpha-1}\,\gamma(1-\alpha,\lambda).$
Thus Assumption (A4) holds.
(5) Since $|\sin x|\leq 1\leq 1+|x|+\mathcal{W}_{2}(\mu,\delta_{0})$ for all $x,\mu$, it’s easy to see that Assumption (A5) holds.
(6) Since $b(x,\mu)=0.1(x-x^{3}+m(\mu))$,
$\langle x,b(x,\mu)\rangle=0.1\big(x^{2}-x^{4}+xm(\mu)\big)\leq 0.1x^{2}+0.1|x||m(\mu)|\leq 0.1x^{2}+0.05x^{2}+0.05\mathcal{W}_{2}^{2}(\mu,\delta_{0}).$
Next,
$\frac{p_{0}-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})$ $=\frac{p_{0}-1}{2}\cdot 0.0125x^{2}.$
Using $|c(x,\mu)|^{2}=\sin^{2}(x)\leq 1$ and $L_{3}=1$, we have
$\langle x,b(x,\mu)\rangle+\frac{p_{0}-1}{2}(|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2})+\frac{1}{2L_{3}}|c(x,\mu)|^{2}\int_{\mathbb{R}_{0}}|z|\Big((1+L_{3}|z|)^{p_{0}-1}-1\Big)\nu(\mathrm{d}z)$
$\quad\leq\Big(0.15+\frac{p_{0}-1}{2}\cdot 0.0125\Big)x^{2}+0.05\mathcal{W}_{2}^{2}(\mu,\delta_{0})+\tfrac{1}{2}\int_{\mathbb{R}_{0}}|z|\Big((1+|z|)^{p_{0}-1}-1\Big)\cdot C_{\alpha,\lambda}\,|z|^{-1-\alpha}e^{-\lambda|z|}\,\mathrm{d}z.$
Let
$I:=\tfrac{1}{2}\int_{\mathbb{R}_{0}}|z|\Big((1+|z|)^{p_{0}-1}-1\Big)\nu(\mathrm{d}z).$
Setting $t=\lambda|z|$, we obtain
$I$ $=C_{\alpha,\lambda}\int_{0}^{\infty}\Big(\frac{t}{\lambda}\Big)^{-\alpha}\Big(\bigl(1+\tfrac{t}{\lambda}\bigr)^{p_{0}-1}-1\Big)e^{-t}\,\frac{\mathrm{d}t}{\lambda}$
$=C_{\alpha,\lambda}\,\lambda^{\alpha-1}\int_{0}^{\infty}t^{-\alpha}\Big(\bigl(1+\tfrac{t}{\lambda}\bigr)^{p_{0}-1}-1\Big)e^{-t}\,\mathrm{d}t.$
As $t\to 0$, the term $\bigl(1+\tfrac{t}{\lambda}\bigr)^{p_{0}-1}-1\sim(p_{0}-1)\tfrac{t}{\lambda}$, so it is integrable if $\alpha<2$. As $t\to\infty$, exponential damping $e^{-t}$ ensures integrability for all $p_{0}$. Thus the integral $I$ is finite for all $p_{0}\in\mathbb{R}$ and $\alpha\in(0,1)$. Hence we may choose
$\gamma_{1}:=0.15+\frac{p_{0}-1}{2}\cdot 0.0125,\qquad\gamma_{2}:=0.05,\qquad\eta:=\tfrac{1}{2}\int_{\mathbb{R}_{0}}|z|\Big((1+|z|)^{p_{0}-1}-1\Big)\nu(\mathrm{d}z).$
This proves Assumption (A6).
(7) Note that
$|\sigma(x)-\sigma(\overline{x})|^{2}+|\sigma^{0}(x)-\sigma^{0}(\overline{x})|^{2}+|c(x)-c(\overline{x})|^{2}\int_{\mathbb{R}_{0}}|z|^{2}\,\nu(\mathrm{d}z)$
$\quad=0.01|x-\overline{x}|^{2}+0.0025|x-\overline{x}|^{2}+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)|\sin x-\sin\overline{x}|^{2}$
$\quad\leq(0.0125+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha))|x-\overline{x}|^{2}.$
Hence choose $\tilde{L}_{1}:=0.0125+2C_{\alpha,\lambda}\,\lambda^{\alpha-2}\,\Gamma(2-\alpha)$ and $\tilde{L}_{2}:=0$. This verifies Assumption (A7).
Remark 2.2.
We point out that assumptions (A1), (A2), (A4), (A6) and (A7) does not hold for an isotropic $\alpha$-stable Lévy measure with index $\alpha\in(0,2)$, and the measure $\nu(\mathrm{d}z)=c_{d,\alpha}\,|z|^{-d-\alpha}\,\mathrm{d}z.$ In fact
$\int_{|z|>1}|z|^{2}\,\nu(\mathrm{d}z)=c_{d,\alpha}\int_{|z|>1}|z|^{2-d-\alpha}\,\mathrm{d}z.$
Using spherical coordinates,
$\int_{|z|>1}|z|^{2-d-\alpha}\,\mathrm{d}z=\int_{1}^{\infty}\int_{S^{d-1}}r^{2-d-\alpha}\,r^{d-1}\,\mathrm{d}\sigma(\theta)\,\mathrm{d}r=S_{d-1}\int_{1}^{\infty}r^{1-\alpha}\,\mathrm{d}r,$
where $S_{d-1}=2\pi^{d/2}/\Gamma(d/2)$ is the surface area of the (d1)(d-1)-sphere.
The radial integral is
$\int_{1}^{\infty}r^{1-\alpha}\,\mathrm{d}r=\frac{1}{2-\alpha}\left(r^{2-\alpha}\right)\Big|_{1}^{\infty}.$
For $\alpha\in(0,2)$ we have $2-\alpha>0$, hence the limit as $r\to\infty$ diverges. Thus
$\int_{|z|>1}|z|^{2}\,\nu(\mathrm{d}z)=\infty.$
Therefore the isotropic $\alpha$-stable Lévy measure does not admit a finite second moment.
@@
To facilitate readability, we divide the proof into the following steps.